Verify that the two given functions are inverses of each other

f(x)=x3+1andg(x)=√3 x−1


1
Expert's answer
2021-11-29T13:41:46-0500

"f(x)=x^3+1"

Domain: "(-\infin, \infin)"

Range: "(-\infin, \infin)"


"f(g(x))=(\sqrt[3]{x-1})^3+1=x-1+1=x, x\in \R"

"g(f(x))=\sqrt[3]{(x^3+1)-1}=\sqrt[3]{x^3}=x, x\in \R"

Therefore the functions"f" and "g" are inverse functions.